# Self-trapping and skin solitons in two-dimensional non-Hermitian lattices

### v1

@@ Abstract
In the context of non-Hermitian photonics, we systematically investigate the spectral and dynamical aspects of two-dimensional nonlinear Hatano-Nelson lattices. In particular, we examine the interplay between self-trapping, due to nonlinearity, and propagation due to the non-Hermitian skin effect. These antagonistic effects give rise to amplitude thresholds - above which self-trapping in a single-channel dominates - that depend highly on the position of initial excitation and the degree of coupling asymmetry. Additionally, we identify two-dimensional skin soliton solutions and characterize their power thresholds and spatial profiles.
preprint: APS/123-QED
I INTRODUCTION
@@
In recent years, photonics has emerged as a versatile platform for exploring non-Hermitian physics, enabling the precise engineering of gain and loss distributions in optical systems [1]. Unlike quantum or condensed matter systems, where the experimental control of non-Hermiticity remains challenging, photonic systems offer a practical and accessible setup [2]. This has fueled rapid progress in the study of non-Hermitian wave dynamics and the demonstration of novel phenomena, such as exceptional points [3, 4], parity-time $(\mathcal{PT})$ symmetry [5, 6, 7, 8, 9, 10, 11, 12, 13], and the Non-Hermitian Skin Effect (NHSE) [14, 15, 16, 17], all of which are now of crucial importance and relevance to the fields of topological and integrated photonics.
Among the plethora of systems studied in non-Hermitian physics, the Hatano-Nelson (HN) model offers a prototypical lattice, where non-Hermiticity is realized through asymmetric nearest-neighbor couplings [18, 19]. An intriguing feature of this model is the NHSE, i.e., the exponential localization of eigenmodes at the boundary of the system under open boundary conditions (OBC). Initially proposed to describe localization-delocalization transitions in condensed matter systems, the HN model remained for years a purely theoretical concept in mathematical physics [20]. However, recent advances in optics have enabled its experimental realization [21, 22], opening new possibilities for practical applications in photonic cavities and lasers [23, 24, 25, 26], as well as, ultracold atoms [27].
@@
The extension of the NHSE to two-dimensional (2D) systems reveals even richer localization features, that depend on the direction and strength of coupling asymmetry [28, 29, 30]. When the coupling asymmetry is along a single direction, NHSE manifests as side-localized modes; however, when asymmetric couplings are present in both directions, the eigenmodes exhibit significant spatial asymmetry as they become localized at specific lattice corners. Notably, most studies related to higher-dimensional lattices have focused on the topological nature of the NHSE, exploring bulk–boundary correspondence and redefining topological invariants, [31, 32, 33, 34, 35, 36, 37], while much less attention has been devoted to transport features.
@@
On the other hand, nonlinearity, an inherent property of optical systems, gives rise to a wide range of intriguing phenomena, including self-trapping, modulational instability, and soliton formation [38]. When combined with non-Hermiticity, the nonlinear phenomena are significantly modified, leading to novel phenomena and the emergence of localized states [39, 40, 41]. In one-dimensional (1D) HN systems, nonlinearity has primarily been explored in the context of topological edge states and single-mode lasing [42, 43, 44, 45]. Recent studies have also examined the effect of Kerr nonlinearity in 1D HN lattices, focusing on soliton formation [46, 47, 48, 49, 50]. Notably, it was recently shown, both theoretically and experimentally, that the interplay between Kerr nonlinearity and asymmetric couplings can lead to the formation of nonlinear skin solitons [46, 47], characterized by position-dependent power thresholds and highly asymmetric spatial profiles. Nevertheless, the interplay of nonlinearity and non-Hermiticity in higher-dimensional systems, such as 2D HN lattices, remains largely unexplored.
@@
In this work, we study the dynamics in 2D HN lattices with Kerr nonlinearity, focusing on the antagonism between self-trapping, and propagation due to asymmetric couplings. In the extensively studied Hermitian case, where the evolution is described by the discrete nonlinear Schrödinger equation (DNLSE), self-trapping occurs when the amplitude of excitation exceeds a critical threshold, suppressing thus discrete diffraction. This phenomenon has been widely explored in both 1D and 2D systems within the framework of lattice solitons [51, 52, 53, 54, 55, 56, 57]. In particular, we examine how the degree of non-Hermiticity and the location of the initial excitation influence the amplitude thresholds for self-trapping. Our study reveals that these thresholds are highly position-dependent, with relatively low thresholds near the lattice corners toward which the couplings are stronger, and significantly higher thresholds — or even impossible self-trapping — in the bulk of the system, under strong non-Hermiticity. Additionally, we identify two-dimensional skin soliton solutions and characterize their properties through their corresponding power-eigenvalue diagrams. Our results show that these skin solitons exhibit spatial asymmetry, and their power thresholds increase monotonically as the coupling asymmetry becomes stronger. We believe that our study may be highly relevant to the emerging direction of nonlinear and non-Hermitian photonics, where the impact of coupling asymmetry to the lasing skin modes is under investigation.
II TWO-DIMENSIONAL NONLINEAR HATANO-NELSON LATTICE
Figure 1: Schematic representation of a two-dimensional Hatano-Nelson lattice. The arrows represent the asymmetric couplings: $e^{h_{x}}$ and $e^{-h_{x}}$ along the $n_{x}$-direction, and $e^{h_{y}}$ and $e^{-h_{y}}$ along the $n_{y}$-direction. The figure identifies the four geometrically distinct corners of the lattice, labeled as (A), (Bx), (By), and (C). The spatial distribution of four indicative linear eigenmodes is included as inset.
Figure 2: Propagation in a 2D HN lattice for the linear (top row) and nonlinear (bottom row) cases, with non-Hermiticity parameter $h=0.2$, and initial condition $\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},13}\delta_{n_{y},13}$. (a)-(c) Normalized complex amplitude $|\psi_{n_{x},n_{y}}(z)|$ for the linear case at propagation distances (a) $z=1$, (b) $z=3.5$, and (c) $z=6$. (d)-(f) Normalized complex amplitude $|\psi_{n_{x},n_{y}}(z)|$ for the nonlinear case and input amplitude $A=4$, for the same propagation distances as before.
II.1 Coupled mode equations
@@
$\begin{split}i\frac{\partial\psi_{n_{x},n_{y}}}{\partial z}+e^{-h_{x}}\psi_{n_{x}+1,n_{y}}+e^{h_{x}}\psi_{n_{x}-1,n_{y}}+e^{-h_{y}}\psi_{n_{x},n_{y}+1}\\ +e^{h_{x}}\psi_{n_{x},n_{y}-1}+g|\psi_{n_{x},n_{y}}|^{2}\psi_{n_{x},n_{y}}=0\end{split}$ (1)
@@
where $h_{x},h_{y}\in\mathbb{R}^{+}$ are the non-Hermiticity parameters, in the horizontal ($n_{x}$) and vertical ($n_{y}$) directions, respectively. The quantity $\psi_{n_{x},n_{y}}(z)$ denotes the complex amplitude of the electric field envelope at site $(n_{x},n_{y})$ at propagation distance $z$. The parameter $g$ determines whether the lattice is linear ($g=0$) or exhibits Kerr-type nonlinearity ($g=1$). Throughout this work, we consistently consider OBC in both lattice directions, i.e., $\psi_{0,n_{y}}=\psi_{N+1,n_{y}}=\psi_{n_{x},0}=\psi_{n_{x},N+1}=0$.
@@
The geometry of such a square lattice, schematically depicted in Fig. 1, implies the existence of four geometrically distinct corners. These corners are labeled as follows: corner (A), located in the direction of stronger coupling along both $n_{x}$ and $n_{y}$, where the linear skin modes are localized; corner (C), located in the direction of weaker coupling along both $n_{x}$ and $n_{y}$; and corners (Bx)/(By), located in the directions of stronger coupling along $n_{x}$/$n_{y}$ but weaker coupling along $n_{y}$/$n_{x}$. In the special case where $h_{x}=h_{y}$, corners Bx and By are equivalent by symmetry, and we refer to them in such case as corners (B). We consistently follow this notation throughout this work.
At this point it is important to mention, that this model can be mapped via the imaginary gauge transformation $\alpha_{n_{x},n_{y}}\equiv\psi_{n_{x},n_{y}}e^{-(h_{x}n_{x}+h_{y}n_{y})}$, to a Hermitian tight-binding model, which, in the nonlinear case ($g=1$), is characterized by site-dependent nonlinearity. The complex amplitude evolution in such a lattice, is governed by the following discrete nonlinear Schrödinger (DNLSE) type of equation:
$\begin{split}i\frac{\partial\alpha_{n_{x},n_{y}}}{\partial z}+\alpha_{n_{x}+1,n_{y}}+\alpha_{n_{x}-1,n_{y}}+\alpha_{n_{x},n_{y}+1}\\ +\alpha_{n_{x},n_{y}-1}+ge^{2(h_{x}n_{x}+h_{y}n_{y})}|\alpha_{n_{x},n_{y}}|^{2}\alpha_{n_{x},n_{y}}=0.\end{split}$ (2)
II.2 Impact of nonlinearity on propagation dynamics
In this section, we investigate how the Kerr nonlinearity affects the dynamics in the 2D NLHN lattice. In all numerical results presented in this study, we consider a lattice consisting of $N\times N=25\times 25$ waveguides and equal non-Hermiticity parameters for the two directions, i.e., $h_{x}=h_{y}\equiv h$.
@@ II.2 Impact of nonlinearity on propagation dynamics
First, as shown in the top row of Fig. 2, in the absence of Kerr nonlinearity ($g=0$), the NHSE dictates the dynamics. Specifically, for a non-Hermiticity parameter $h=0.2$ and an initial condition $\psi_{n_{x},n_{y}}(z=0)=A\,\delta_{n_{x},13}\delta_{n_{y},13},$ where $A$ is the excitation amplitude, the wave function propagates towards the corner (A) of the lattice, as is physically expected. In the linear regime, the diffraction is, by default, independent of the excitation amplitude $A$, unlike the nonlinear case, where its value plays a critical role. In particular, as illustrated in the bottom row of Fig. 2, for the same initial condition with amplitude $A=4$, the wavefunction remains self-trapped at its initial position for short propagation distances ($z<2$). Then, for $z\gtrsim 2$, the wavefunction splits into two parts: one remains localized at the site of the initial condition, while the other propagates toward corner (A). This behavior clearly indicates an antagonism between self-trapping caused by nonlinearity, and the NHSE, caused by the asymmetric couplings.
Figure 3: Evolution of mean positions and uncertainties in a 2D HN lattice with non-Hermiticity parameter $h=0.2$, for the initial condition $\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},13}\delta_{n_{y},13}$. (a) Mean positions $\langle n_{x}\rangle$ and $\langle n_{y}\rangle$ and (b) position uncertainties $\Delta n_{x}$ and $\Delta n_{y}$ for the linear case, blue line) and nonlinear cases with excitation amplitudes $A=1$ (red line), $A=4$ (yellow line), and $A=6$ (magenta line).
Figure 4: Role of $h$ on self-trapping, for fixed position of the excitation channel at the center of the lattice, $\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},13}\delta_{n_{y},13}$. Normalized amplitude $|\psi_{n_{x},n_{y}}(z_{\text{max}})|$ for different values of $A$ (left panels) and $IPR$ versus $A$ (right panels). Red dots correspond to the values of $A$ in the left panels. Two cases are considered: (a) $h=0$ and (b) $h=0.2$.
@@
In order to further understand such an effect, we also study the evolution of the wavefunction’s mean positions and uncertainties along the $n_{x}$ and $n_{y}$ directions. These are defined by the following general expressions:
@@
Our results are presented in Fig.3. In particular, in Fig. 3(a), we show that, in the linear regime, the wavefunction’s mean positions, $\langle n_{x}\rangle=\langle n_{y}\rangle$, increase almost linearly with the propagation distance $z$, after an initial acceleration [36], and finally the wavepacket reaches the corner C. To the contrary, when Kerr nonlinearity is introduced, propagation is significantly hindered. Specifically, as the excitation amplitude $A$ increases, the mean positions $\langle n_{x}\rangle=\langle n_{y}\rangle$ have a progressively delayed increase and their value never reaches the corresponding for the linear regime, indicating a higher tendency for self-trapping. Moreover, the nonlinear regime is characterized by increased position uncertainties $\Delta n_{x}=\Delta n_{y}$, as depicted in Fig. 3(b). This increase reflects the previously discussed splitting of the wavefunction into two parts (bottom row of Fig. 2). Such behavior emerges due to the interplay between Kerr nonlinearity and the NHSE, which will be studied more systematically in the following section.
III ANTAGONISM BETWEEN SKIN EFFECT AND SELF-TRAPPING
Figure 5: Role of the location of the excitation channel on self-trapping, for fixed $h=0.4$. Amplitude profiles $|\psi_{n_{x},n_{y}}(z_{\text{max}})|$ for various values of $A$ (top-row) under single-channel excitation at different lattice sites (a) at the center, $(n_{x_{0}},n_{y_{0}})=(13,13)$; (b) near corner (A), $(n_{x_{0}},n_{y_{0}})=(23,23)$; and (c) near corner (B), $(n_{x_{0}},n_{y_{0}})=(23,3)$. $IPR$ versus $A$ is shown in bottom-row for all cases. Red dots indicate the values of $IPR$ for the corresponding values of $A$ in top row.
As it was discussed in the previous paragraphs, the skin effect and nonlinearity have opposite tendencies regarding localization, which leads to a dynamic interplay. Therefore, our goal in this section is to demonstrate, through pertinent examples, that the dynamics in a 2D NLHN lattice, as described by Eq. (1) with $g=1$, depend highly on the position of the initial excitation, due to the system’s non-Hermiticity. Specifically, we will consider the single-site initial condition:
@@ III ANTAGONISM BETWEEN SKIN EFFECT AND SELF-TRAPPING
First, we focus on the case of initial excitation at the center of the lattice, i.e., $(n_{x_{0}},n_{y_{0}})=(13,13)$ for Hermitian and non-Hermitian lattices and our results are shown in Fig. 4. In particular, in Fig. 4(a) for the Hermitian case ($h=0$), the wavefunction at a fixed propagation distance $z_{\text{max}}=5$, $|\psi_{n_{x},n_{y}}(z_{\text{max}})|$, is extended for low values of the excitation amplitude $A$. When $A$ exceeds a particular value, the effects of Kerr nonlinearity become significant, leading to almost total self-trapping of the wavefunction at the position of initial excitation. The transition between delocalization and self-trapping can be quantified by the increase in the output (i.e., at $z=z_{\text{max}}$) value of the inverse participation ratio ($IPR$). This metric is widely used to indicate localization and, for a given two-dimensional wavefunction $\psi_{n_{x},n_{y}}$, is defined as:

### v2

@@ Abstract
Two-dimensional non-Hermitian photonic lattices with asymmetric couplings offer rich possibilities for controlling wave localization, through the emergence of the non-Hermitian skin effect at lattice corners or sides. Incorporating optical nonlinearity fundamentally alters these boundary-localization characteristics. Here we show that in a two-dimensional Hatano–Nelson lattice with Kerr nonlinearity, the interplay between self-trapping and directional propagation leads to position-dependent amplitude thresholds. Single-site excitations having above a critical amplitude become confined to their initial position, with lower thresholds near the position where the linear eigenmodes are localized and higher thresholds within the lattice’s bulk. Additionally, we study the differences of this dynamical interplay, for wider initial excitations, between the focusing and defocusing Kerr-nonlinearity regimes. Lastly, we identify skin soliton solutions in a variety of two-dimensional lattice geometries featuring coupling asymmetry.
Introduction
@@
In recent years, photonics has emerged as a versatile platform for exploring non-Hermitian physics, enabling the precise engineering of gain and loss distributions in optical systems [1]. Unlike quantum or condensed-matter systems, where the experimental control of non-Hermiticity remains challenging, photonic systems offer a practical and accessible setup [2]. This has fueled rapid progress in the study of non-Hermitian wave dynamics and the demonstration of novel phenomena, such as exceptional points [3, 4], parity-time $(\mathcal{PT})$ symmetry [5, 6, 7, 8, 9, 10, 11, 12, 13, 14], and the Non-Hermitian Skin Effect (NHSE) [15, 16, 17, 18], among other [19, 20, 21, 22, 23], all of which are now of crucial importance and relevance to the fields of topological and integrated photonics.
@@
Among the plethora of systems studied in non-Hermitian physics, the Hatano–Nelson (HN) model offers a prototypical lattice where non-Hermiticity is realized through asymmetric nearest-neighbor couplings [24, 25]. An intriguing feature of this model is the NHSE, i.e., the exponential localization of eigenmodes at the boundary of the system under open boundary conditions (OBC). Initially proposed to describe localization–delocalization transitions in solid-state systems, the HN model remained for years a purely theoretical concept in mathematical physics [26]. However, recent advances in optics have enabled its experimental realization [27, 28], opening new possibilities for practical applications in photonic cavities and lasers [29, 30, 31, 32, 33] as well as ultracold atoms [34]. The extension of the NHSE to two-dimensional (2D) systems reveals even richer localization features that depend on the direction and strength of coupling asymmetry [35, 36, 37, 38, 39]. When the coupling asymmetry is along a single direction, the NHSE manifests as side-localized modes; however, when asymmetric couplings are present in both directions, the eigenmodes exhibit significant spatial asymmetry as they become localized at specific lattice corners. Notably, most studies related to higher-dimensional lattices have focused on the topological nature of the NHSE, exploring bulk–boundary correspondence and redefining topological invariants [40, 41, 42, 43, 44, 45, 46], while much less attention has been devoted to transport features [38].
@@
On the other hand, nonlinearity, an inherent property of optical systems, gives rise to a wide range of intriguing phenomena, including self-trapping, modulational instability, and soliton formation [47]. When combined with non-Hermiticity, the nonlinear phenomena are significantly modified, leading to novel effects and the emergence of localized states [48, 49, 50]. In one-dimensional (1D) HN systems, nonlinearity has primarily been explored in the context of topological edge states and single-mode lasing [51, 52, 53, 54]. Recent studies have also examined the effect of Kerr nonlinearity in 1D HN lattices, focusing on soliton formation [55, 56, 58, 59, 61, 60]. Notably, it was recently shown, both theoretically and experimentally, that the interplay between Kerr nonlinearity and asymmetric couplings can lead to the formation of nonlinear skin solitons [55, 56, 57], characterized by position-dependent power thresholds and highly asymmetric spatial profiles. Nevertheless, the interplay of nonlinearity and non-Hermiticity in higher-dimensional systems, such as 2D HN lattices, remains largely unexplored.
@@
In this work, we study the dynamics in 2D HN lattices with Kerr nonlinearity, focusing on the antagonism between self-trapping and propagation due to asymmetric couplings. In the extensively studied Hermitian case, where the evolution is described by the discrete nonlinear Schrödinger equation (DNLSE), self-trapping of a localized excitation to its initial position occurs when the excitation’s amplitude exceeds a critical threshold, thus suppressing discrete diffraction. This phenomenon has been widely explored in both 1D and 2D systems within the framework of lattice solitons [62, 63, 64, 65, 66, 67, 68]. In particular, here we examine how the degree of non-Hermiticity and the location of the initial single-site excitation influence the amplitude thresholds for self-trapping. Our study reveals that these thresholds are highly position-dependent, with relatively low thresholds near the lattice corners toward which the couplings are stronger, and significantly higher thresholds—or even impossible self-trapping—in the bulk of the system under strong non-Hermiticity. Furthermore, we consider wider Gaussian initial excitations and investigate how the interplay between Kerr nonlinearity and non-Hermiticity depends on whether the nonlinearity is focusing or defocusing. Additionally, we identify two-dimensional skin soliton solutions and characterize their properties through their corresponding power–eigenvalue diagrams. Our results show that these skin solitons exhibit spatial asymmetry, and their power thresholds increase monotonically as the coupling asymmetry becomes stronger. Finally, we report the existence of skin solitons in a variety of more complex two-dimensional nonlinear non-Hermitian lattices, beyond the 2D HN lattice.
Results
Two-dimensional nonlinear Hatano-Nelson lattice
Coupled-mode equations and conservation laws
@@ Results
$\begin{split}i\frac{\partial\psi_{n_{x},n_{y}}}{\partial z}+e^{-h_{x}}\psi_{n_{x}+1,n_{y}}+e^{h_{x}}\psi_{n_{x}-1,n_{y}}+e^{-h_{y}}\psi_{n_{x},n_{y}+1}+e^{h_{y}}\psi_{n_{x},n_{y}-1}+g|\psi_{n_{x},n_{y}}|^{2}\psi_{n_{x},n_{y}}=0\end{split}$ (1)
@@
where $h_{x},h_{y}\in\mathbb{R}^{+}$ are the non-Hermiticity parameters, in the horizontal ($n_{x}$) and vertical ($n_{y}$) directions, respectively. The quantity $\psi_{n_{x},n_{y}}(z)$ denotes the complex amplitude of the electric field’s envelope at site $(n_{x},n_{y})$ at propagation distance $z$. The parameter $g$ determines whether the lattice is linear ($g=0$) or exhibits Kerr-type nonlinearity ($|g|=1$). Throughout this work, we consistently consider OBC in both lattice directions, i.e., $\psi_{0,n_{y}}=\psi_{N+1,n_{y}}=\psi_{n_{x},0}=\psi_{n_{x},N+1}=0$.
Figure 1: Schematic representation of a two-dimensional Hatano-Nelson lattice. The arrows represent the asymmetric couplings: $e^{h_{x}}$ and $e^{-h_{x}}$ along the $n_{x}$-direction, and $e^{h_{y}}$ and $e^{-h_{y}}$ along the $n_{y}$-direction. The figure identifies the four geometrically distinct corners of the lattice, labeled as (A), (Bx), (By), and (C). The spatial distribution of four indicative linear eigenmodes, for a case with $h_{x}=h_{y}$, is included as inset.
@@
The geometry of such a square lattice, schematically depicted in Fig. 1, implies the existence of four geometrically distinct corners. These corners are labeled as follows: corner (A), located in the direction of stronger coupling along both $n_{x}$ and $n_{y}$, where the linear skin modes are localized; corner (C), located in the direction of weaker coupling along both $n_{x}$ and $n_{y}$; and corners (Bx)/(By), located in the directions of stronger coupling along $n_{x}$$n_{y}$ but weaker coupling along $n_{y}$$n_{x}$. In the special case where $h_{x}=h_{y}$, corners Bx and By are equivalent by symmetry, and we refer to them in such case as corners (B). We consistently follow this notation throughout this work.
Although the 2D NLHN is non-Hermitian and thus corresponds to non-conservative dynamics, it can be mapped via the imaginary gauge transformation $\alpha_{n_{x},n_{y}}\equiv\psi_{n_{x},n_{y}}e^{-(h_{x}n_{x}+h_{y}n_{y})}$, to a Hermitian tight-binding model, which, in the nonlinear case ($|g|=1$), is characterized by site-dependent nonlinearity. As we will see, this allow us to derive conservation laws for the system under discussion.
Specifically, the evolution of the complex amplitude $\alpha_{n_{x},n_{y}}(z)$ in such a lattice is governed by the discrete nonlinear Schrödinger (DNLSE)-type equation:
$\begin{split}i\frac{\partial\alpha_{n_{x},n_{y}}}{\partial z}+\alpha_{n_{x}+1,n_{y}}+\alpha_{n_{x}-1,n_{y}}+\alpha_{n_{x},n_{y}+1}+\alpha_{n_{x},n_{y}-1}+ge^{2(h_{x}n_{x}+h_{y}n_{y})}|\alpha_{n_{x},n_{y}}|^{2}\alpha_{n_{x},n_{y}}=0.\end{split}$ (2)
As detailed in the Supplemental Material (SM), the quantity
$\mathcal{P}(z)\equiv\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\alpha_{n_{x},n_{y}}(z)|^{2}$ (3)
which represents the optical power and is known to be conserved in the standard DNLSE, remains conserved even under this particular site-dependent nonlinearity. This contrasts with the Hatano-Nelson system of Eq. (1), in which the optical power $\mathcal{P}_{\text{HN}}(z)=\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\psi_{n_{x},n_{y}}(z)|^{2}$ is not conserved along propagation. Nonetheless, the connection between the systems described by Eq. (1) and Eq. (2) through the transformation $\alpha_{n_{x},n_{y}}\equiv\psi_{n_{x},n_{y}}e^{-(h_{x}n_{x}+h_{y}n_{y})},$ along with the power conservation of the latter system, implies that in the 2D NLHN lattice [Eq. (1)], the quantity
$\tilde{\mathcal{P}}=\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\psi_{n_{x},n_{y}}|^{2}e^{-2(h_{x}n_{x}+h_{y}n_{y})}$ (4)
which we refer to as the pseudopower, is constant during the dynamics. Moreover, another invariant quantity of the system described by Eq.(2) is the Hamiltonian,
$\begin{split}\mathcal{H}\equiv&\sum_{n_{x}=1}^{N-1}\sum_{n_{y}=1}^{N-1}\big(\alpha_{n_{x},n_{y}}\alpha^{*}_{n_{x}+1,n_{y}}+\alpha_{n_{x},n_{y}}\alpha^{*}_{n_{x},n_{y}+1}+\text{c.c.}\big)+\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}ge^{2(h_{x}n_{x}+h_{y}n_{y})}\frac{|\alpha_{n_{x},n_{y}}|^{4}}{2},\end{split}$ (5)
as shown in the SM. Consequently, a second conserved quantity for the 2D NLHN lattice is the pseudohamiltonian
$\begin{split}\tilde{\mathcal{H}}=&\sum_{n_{x}=1}^{N-1}\sum_{n_{y}=1}^{N-1}e^{-2(h_{x}n_{x}+h_{y}n_{y})}\Big(2e^{-h_{x}}\,\text{Re}(\psi_{n_{x},n_{y}}^{*}\psi_{n_{x}+1,n_{y}})+2e^{-h_{y}}\,\text{Re}(\psi_{n_{x},n_{y}}^{*}\psi_{n_{x},n_{y}+1})\Big)+\sum_{n_{x}=1}^{N-1}\sum_{n_{y}=1}^{N-1}ge^{-2(h_{x}n_{x}+h_{y}n_{y})}\frac{|\psi_{n_{x},n_{y}}|^{4}}{2}\end{split}$ (6)
Within the context of our discussion, the existence of these two conservation laws provides a direct means of validating the accuracy of our numerical results, which will be presented in the following subsections. This is very important, given the high degree of non-normality and numerical instability due to coupling asymmetry, that are inherent in Hatano-Nelson lattices [69].
Impact of nonlinearity on propagation dynamics
In this paragraph, we investigate how the Kerr nonlinearity affects the dynamics in the 2D NLHN lattice. Unless stated differently, in the numerical results presented in this study, we consider a lattice consisting of $N\times N=25\times 25$ waveguides and equal non-Hermiticity parameters for the two directions, i.e., $h_{x}=h_{y}\equiv h$. Complementary results for the case of $h_{x}\neq 0$ and $h_{y}=0$ are discussed in the SM. In particular, we will consider the scenario of a localized, single-site initial excitation. As analytically proven in the SM, for such an initial condition, the dynamics of the magnitude of the wavefunction $|\psi_{n_{x},n_{y}}(z)|$ remain identical regardless of whether the Kerr nonlinearity is focusing ($g=1$) or defocusing (g=1g=-1). Relevant examples of propagation dynamics with initial excitations occupying multiple channels are discussed in a separate subsequent subsection.
Figure 2: Propagation in a 2D HN lattice for the linear (top row) and Kerr-nonlinear (bottom row) cases, with non-Hermiticity parameter $h=0.2$ and initial condition $\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},13}\delta_{n_{y},13}$. (a)–(c) Normalized complex amplitude $|\psi_{n_{x},n_{y}}(z)|$ for the linear case at propagation distances (a) $z=1$, (b) $z=3.5$, and (c) $z=6$. (e)–(g) Normalized complex amplitude $|\psi_{n_{x},n_{y}}(z)|$ for the nonlinear case with input amplitude $A=4$, shown at the same propagation distances as in the top row. Panels (d) and (h) show the evolution of the optical power $\mathcal{P}_{\text{HN}}$ (red lines) and the optical intensity at the site of initial excitation $|\psi_{13,13}|^{2}$ (blue lines) for the linear and nonlinear cases, respectively.
@@
First, as shown in the top row of Fig. 2, in the absence of Kerr nonlinearity ($g=0$), the NHSE dictates the dynamics. Specifically, for a non-Hermiticity parameter $h=0.2$ and an initial condition $\psi_{n_{x},n_{y}}(z=0)=A\,\delta_{n_{x},13}\delta_{n_{y},13}$, where $A$ is the excitation amplitude, the wavefunction propagates toward corner (A) of the lattice, as physically expected. As clearly demonstrated in Fig. 2(d), while the wavefunction shifts toward corner (A), the total optical power $\mathcal{P}_{\text{HN}}=\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\psi_{n_{x},n_{y}}(z)|^{2}$ increases, consistent with the conservation of the pseudopower $\tilde{\mathcal{P}}=\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\psi_{n_{x},n_{y}}(z)|^{2}e^{-2(h_{x}n_{x}+h_{y}n_{y})}$ , since the wavefunction is on average distributed on lattice sites with larger $n_{x}$, $n_{y}$. Additionally, the optical intensity at the initially excited site, $|\psi_{13,13}|^{2}$, vanishes for $z>1$. Of course, in the linear regime, diffraction is independent of the excitation amplitude $A$, unlike the nonlinear case ($|g|=1$), where its value plays a crucial role. In particular, as illustrated in the bottom row of Fig. 2, for the same initial condition, with amplitude $A=4$, the wavefunction remains self-trapped at the initial-excitation site for short propagation distances ($z<2$). For $z\gtrsim 2$, it splits into two distinct parts: one remains localized at the initial position, while the other propagates toward corner (A). This occurs because, in the Kerr nonlinear case, the optical intensity $|\psi_{13,13}|^{2}$ exhibits oscillations for $z<2$ [Fig. 2(h)] and eventually stabilizes at a slightly reduced value compared to its initial one. The separated portion of optical power propagates toward the preferential corner and becomes amplified due to the conservation of pseudopower $\mathcal{\tilde{P}}$, as in the linear regime. This process results in wavefunction splitting: part of the initial power, slightly less than the original, remains localized near the initial-excitation site, while the rest is amplified as it propagates toward the preferential corner. This behavior clearly indicates an antagonism between self-trapping caused by nonlinearity, and the NHSE, induced by the asymmetric couplings.
@@
In order to better understand this interplay, we study the evolution of the wavefunction’s mean positions and uncertainties along the $n_{x}$ and $n_{y}$ directions. These are defined by the following general expressions:
@@
Our corresponding results, for the same configuration as in Fig.2, are presented in Fig.3. In particular, in Fig. 3(a), we show that, in the linear regime, the wavefunction’s mean positions, $\langle n_{x}\rangle=\langle n_{y}\rangle$, increase almost linearly with the propagation distance $z$, after an initial acceleration [45], and finally the wavepacket reaches the corner (A). To the contrary, when Kerr nonlinearity is introduced, propagation is significantly hindered. Specifically, as the excitation amplitude $A$ increases, the mean positions $\langle n_{x}\rangle=\langle n_{y}\rangle$ have a progressively delayed increase and their value never reaches the corresponding for the linear regime, indicating a higher tendency for self-trapping. Moreover, the nonlinear regime is characterized by increased position uncertainties $\Delta n_{x}=\Delta n_{y}$, as depicted in Fig. 3(b). This increase reflects the previously discussed splitting of the wavefunction into two parts (bottom row of Fig. 2). Such behavior emerges due to the interplay between Kerr nonlinearity and the NHSE, which will be studied more systematically in the following subsection.
Figure 3: Evolution of mean positions and uncertainties in a 2D HN lattice with non-Hermiticity parameter $h=0.2$, for the initial condition $\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},13}\delta_{n_{y},13}$. (a) Mean positions $\langle n_{x}\rangle$ and $\langle n_{y}\rangle$ and (b) position uncertainties $\Delta n_{x}$ and $\Delta n_{y}$ for the linear case (gray line) and nonlinear cases with excitation amplitudes $A=1$ (red line), $A=4$ (blue line), and $A=6$ (green line).
Antagonism between skin effect and self-trapping
As it was discussed in the previous section, Kerr-nonlinearity and the skin effect have opposite tendencies regarding localization of an initially localized wavepacket, which leads to a dynamic interplay. Therefore, our goal in this section is to demonstrate, through pertinent examples, that the dynamics in a 2D NLHN lattice, as described by Eq. (1) with $|g|=1$, depend highly on the position of the initial single-site excitation, due to the system’s non-Hermiticity. Specifically, we will consider the single-site initial condition:
@@
First, we focus on the case of initial excitation at the center of the lattice, i.e., $n_{x_{0}}=n_{y_{0}}=13$ for Hermitian and non-Hermitian lattices and our results are shown in Fig. 4. In particular, in Fig. 4(a) for the Hermitian case ($h=0$), the wavefunction at a fixed propagation distance $z=5$, $|\psi_{n_{x},n_{y}}(z=5)|$, is extended for low values of the excitation amplitude $A$. When $A$ exceeds a particular value, the effect of Kerr nonlinearity becomes significant, leading to self-trapping of the wavefunction, almost entirely, at the position of initial excitation. The transition between delocalization and self-trapping can be quantified by the increase in the output (i.e., at $z=5$) value of the inverse participation ratio ($IPR$). This metric is widely used to indicate localization and, for a given two-dimensional wavefunction $\psi_{n_{x},n_{y}}$, is defined as:
@@
Figure 4: Comparison of the self-trapping tendency between (a) a Hermitian ($h=0$) and (b) a non-Hermitian ($h=0.2$) lattice, for single-site excitation at the center, i.e., $\psi_{n_{x},n_{y}}(z=0)=A\,\delta_{n_{x},13}\delta_{n_{y},13}$. The normalized amplitudes $|\psi_{n_{x},n_{y}}(z=5)|$ are plotted for pertinent values of AA (left panels) and the IPRIPR for $z=5$ is shown as a function of AA (right panels). Red dots correspond to the AA values used in the left panels.